# ControlNet+++++ +: Improving Conditional Controls with Efficient Consistency Feedback Project Page: liming-ai.github.io/ControlNet_Plus_Plus

### v1

@@ Abstract
To enhance the controllability of text-to-image diffusion models, existing efforts like ControlNet incorporated image-based conditional controls. In this paper, we reveal that existing methods still face significant challenges in generating images that align with the image conditional controls. To this end, we propose ControlNet++, a novel approach that improves controllable generation by explicitly optimizing pixel-level cycle consistency between generated images and conditional controls. Specifically, for an input conditional control, we use a pre-trained discriminative reward model to extract the corresponding condition of the generated images, and then optimize the consistency loss between the input conditional control and extracted condition. A straightforward implementation would be generating images from random noises and then calculating the consistency loss, but such an approach requires storing gradients for multiple sampling timesteps, leading to considerable time and memory costs. To address this, we introduce an efficient reward strategy that deliberately disturbs the input images by adding noise, and then uses the single-step denoised images for reward fine-tuning. This avoids the extensive costs associated with image sampling, allowing for more efficient reward fine-tuning. Extensive experiments show that ControlNet++ significantly improves controllability under various conditional controls. For example, it achieves improvements over ControlNet by 7.9% mIoU, 13.4% SSIM, and 7.6% RMSE, respectively, for segmentation mask, line-art edge, and depth conditions.
@@ 1 Introduction
The emergence and improvements of diffusion models [50, 12, 43], along with the introduction of large-scale image-text datasets [49, 48], has catalyzed significant strides in text-to-image generation. Nonetheless, as the proverb “an image is worth a thousand words” conveys, it’s challenging to depict an image accurately and in detail through language alone, and this dilemma also perplexes existing text-to-image diffusion models [43, 46]. To this end, many studies focus on incorporating conditional controls such as segmentation mask, and depth map into text-to-image diffusion models [63, 30, 62, 37, 22] to facilitate controllable generation. Despite the diversity in these methods, the core objective remains to facilitate more accurate and controllable image generation with explicit image-based conditional controls.
Achieving controllable image generation could involve retraining diffusion models from scratch [43, 37], but this comes with high computational demands and a scarcity of large public datasets [63]. In light of this, a feasible strategy is fine-tuning pre-trained text-to-image models [23, 61] or introducing trainable modules [63, 30, 62] like ControlNet [63]. However, despite these studies have explored the feasibility of controllability [63, 30, 62] in text-to-image diffusion models and expanded various applications [23, 37, 22], a significant gap remains in achieving precise and fine-grained control. As shown in Fig. 1, existing state-of-the-art methods of controllable generation (e.g., ControlNet [63] and T2I-Adapter [30]) still struggle to accurately generate images that are consistent with the input image condition. For example, T2I-Adapter-SDXL consistently produced incorrect wrinkles on the forehead in all generated images, while ControlNet v1.1 introduced many wrong details. Regrettably, current efforts lack specific methods for improving controllability, which impedes progress in this research area.
@@
  • Consistency Feedback: We show that pre-trained discriminative models can serve as powerful visual reward models to improve the controllability of controllable diffusion models in a cycle-consistency manner.
@@ 2.2 Controllable Text-to-Image Diffusion Models
To achieve conditional control in pre-trained text-to-image diffusion models, ControlNet [63] and T2I-Adapter [30] introduce additional trainable modules for guided image generation. Furthermore, recent research employs various prompt engineering [27, 64, 61] and cross-attention constraints [58, 6, 23] for a more regulated generation. Some methods also explore multi-condition or multi-modal generation within a single diffusion model [37, 21, 65] or focus on the instance-based controllable generation [54, 69]. However, despite these methods exploring feasibility and applications, there still lacks a clear approach to enhance controllability under various conditional controls. Furthermore, existing works implicitly learn controllability by the denoising process of diffusion models, while our ControlNet++ achieves this in a cycle-consistency manner, as shown in Fig. 3.
@@ 3.1 Preliminary
where $\epsilon_{\theta}$ refers to the predicted noise at timestep $t$ by U-Net [44] with parameters $\theta$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of posterior Gaussian distribution $p_{\theta}(x_{0})$
@@ 3.2 Reward Controllability with Consistency Feedback
where $\lambda$ is a hyper-parameter to adjust the weight of the reward loss. Through this approach, the consistency loss can guide the diffusion model on how to sample at different timesteps to obtain images more consistent with the input conditional controls, thereby enhancing controllability. Nonetheless, directly applying such reward consistency still poses challenges in efficiency in real-world settings.
@@ 3.3 Efficient Reward Fine-tuning
$x_{0}\approx x^{\prime}_{0}=\frac{x^{\prime}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x^{\prime}_{t},c_{v},c_{t},t-1\right)}{\sqrt{\alpha_{t}}},$ (7)
@@
During the reward fine-tuning phases, we freeze the pre-trained discriminative reward model and text-to-image model, and only update the ControlNet module following its original implementation, which ensures its generative capabilities are not compromised. In our experiments, we observe that using only the reward loss will lead to image distortion, aligning with the conclusions drawn in previous studies [60].
@@ 4.1 Experimental Setup
Given that existing text-image paired datasets for generative models are unable to provide accurate conditional control data pairs [49, 48], such as image-segmentation pairs, we endeavor to select specific datasets for different tasks that can offer more precise image-label data pairs. More specifically, ADE20K [67, 68] and COCOStuff [4] are used for the segmentation mask condition following ControlNet [63]. For the canny edge map, hed edge map, lineart map, and depth map condition, we utilize the MultiGen-20M dataset [37], which is a subset of LAION-Aesthetics [48]. For the datasets without text caption such as ADE20K, we utilize MiniGPT-4 [70] to generate the image caption with the instruction “Please briefly describe this image in one sentence”. The image resolution for both training and inference is 512$\times$512 for all datasets and methods. More details are provided in Sec. 8.
@@
We train ControlNet++ on the training set of each corresponding dataset and evaluate all methods on the validation dataset. All the experiments are evaluated under 512$\times$512 resolution for fair comparison. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated images of diffusion models. For semantic segmentation and depth map controls, we use mIoU and RMSE as evaluation metrics respectively, which is a common practice in related research fields. For the edge task, we use F1-Score for hard edges (canny edge) because it can be regarded as a binary classification problem of 0 (non-edge) and 1 (edge) and has a serious long-tail distribution, following the standard evaluation in edge detection [59]. The threshold used for evaluation is (100, 200) for OpenCV, and (0.1, 0.2) for Kornia implementation. The SSIM metric is used for the soft edges conditional controls (i.e., hed edge & lineart edge) following previous works [65].
@@
Our evaluation primarily focuses on T2I-Adapter [30], ControlNet v1.1 [63], GLIGEN [27], Uni-ControlNet [65], and UniControl [37], as these methods are pioneering in the realm of controllable text-to-image diffusion models and offer public model weights for various image conditions. To ensure fairness of evaluation, all methods use the same image conditions and text prompts. For ControlNet++, we use the UniPC [66] sampler with 20 denoising steps to generate images with the original text prompt following ControlNet v1.1 [63], without any negative prompts. For other methods, we utilized their open-source code to generate images and conducted fair evaluations under the same settings, without changing any configures such as the number of inference steps or denoising sampler. While most methods employ the user-friendly SD1.5 as their text-to-image model for controllable generation, we have observed that recently there are a few models based on SDXL [35]. Therefore, we also report the controllability results for ControlNet-SDXL and T2I-Adapter-SDXL. Please note that ControlNet-SDXL mentioned here is not an officially released model as in ControlNet [63]. More details are provided in the Supplementary Material.
Table 1: Controllability comparison with state-of-the-art methods under different conditional controls and datasets. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. ControlNet++ achieves significant controllability improvements. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images for each result and report the average result.
@@ 4.2 Experimental Results
In this section, we compare our ControlNet++ with several state-of-the-art methods in terms of controllability and image quality under different image-based conditional controls. We employ Stable Diffusion v1.5 (SD1.5) following the previous works [30, 63, 37, 65] for fair comparison.
@@
The experimental results are shown in Tab. 1, which can be summarized as the following observations: (1) Existing methods still underperform in terms of controllability, struggling to achieve precise controlled generation. For instance, current methods (i.e., ControlNet) achieve only a 27.46 mIoU for images generated under the condition of segmentation masks, which is far from its performance on real datasets with a 50.7 mIoU, under the same evaluation from Mask2Former segmentation model [8]. (2) Our ControlNet++ significantly outperforms existing works in terms of controllability across various conditional controls. For example, it achieves 7.6% RMSE improvements against previous state-of-the-art methods for the depth map condition; (3) For controllable diffusion models, the strength of the text-to-image backbone does not affect its controllability. As shown in the table, although SDXL-based [35] ControlNet and T2I-Adapter have better controllability on some specific tasks, the improvement is not large and is not significantly better than the counterparts with SD 1.5 [43].
Qualitative Comparison.
Figs. 5 and 6 provide a qualitative comparison between our ControlNet++ and previous state-of-the-art methods across different conditional controls. When given the same input text prompts and image-based conditional controls, we observe that existing methods often generate areas inconsistent with the image conditions. For instance, in the segmentation mask generation task, other methods often produce extraneous picture frames on the walls,
Figure 6: Visualization comparison results on Line-Art Edge.
resulting in a mismatch between the segmentation masks extracted from the generated image and the input segmentation mask condition. A similar situation occurs under depth conditions, where other methods fail to accurately represent the depth information of different fingers. In contrast, images generated by our ControlNet++ maintain good consistency with the input depth map.
Table 2: FID ($\downarrow$) comparison with state-of-the-art methods under different conditional controls and datasets. All the results are conducted on 512$\times$512 image resolution with Clean-FID implementation [33] for fair comparisons. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images for each result and report the average result to reduce random errors.
@@
Table 3: CLIP-score ($\uparrow$) comparison with state-of-the-art methods under different conditional controls and datasets. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images for each result and report the average result to reduce random errors.
@@
To verify whether improving controllability leads to a decline in image quality, we reported the FID (Fréchet Inception Distance) metrics of different methods under various conditional generation tasks in Tab. 2. We discovered that, compared to existing methods, ControlNet++ generally exhibits superior FID values in most cases, indicating that our approach, while enhancing the controllability of conditional controls, does not result in a decrease in image quality. This can also be observed in Fig. 5. We provide more visual examples in Figures 12,13,14,15,16.
@@
To further validate our improvements in controllability and their impact, we use the generated images along with real human-annotated labels to create a new dataset for training discriminative models from scratch. Please note that the only difference from the original dataset used to train the discriminative model is that we have replaced the images with those generated by the controllable diffusion model while keeping the labels unchanged. If the generative model exhibits good controllability, the quality of the constructed dataset will be higher, thereby enabling the training of a more powerful discriminative model.
@@
In summary: (1) both experiments (training the discriminative model on generated images only or on the real and generated images) show that our method is better than ControlNet, demonstrating better controllability is beneficial for training discriminative models. (2) Our generated image can help significantly boost the discriminative model (e.g., segmentation model) performance when training with real data.
@@ 4.3 Ablation Study
In our experiments, we find that maintaining the original diffusion training process is crucial for preserving the quality and controllability of generated images. Relying solely on pixel-level consistency loss leads to severe image distortion, whereas training the model with both this loss and the diffusion training loss can enhance controllability without affecting image quality. We demonstrate the visualization results in Fig. 8.
Influence of Text Prompt.
Following ControlNet [63], here we discuss how different types of text prompts (No Prompt, Conflicting Prompt, and Perfect Prompt.) affect the final experimental results. As shown in Fig. 9, when the text prompt is empty or there is a semantic conflict with the image conditional control, ControlNet often struggles to generate accurate content. In contrast, our ControlNet++ manages to generate images that comply with the input conditional controls under various text prompt scenarios.
Table 4: Stronger reward model (UperNet) leads to better controllability than the weaker reward model (DeepLabv3). The eval mIoU are measured by Mask2Former with performance of 56.01 mIoU.
Reward Model (RM) RM mIoUnormal-$\uparrow$ Eval mIoUnormal-$\uparrow$
- - 31.38

### v2

@@ Abstract
To enhance the controllability of text-to-image diffusion models, existing efforts like ControlNet incorporated image-based conditional controls. In this paper, we reveal that existing methods still face significant challenges in generating images that align with the image conditional controls. To this end, we propose ControlNet++, a novel approach that improves controllable generation by explicitly optimizing pixel-level cycle consistency between generated images and conditional controls. Specifically, for an input conditional control, we use a pre-trained discriminative reward model to extract the corresponding condition of the generated images, and then optimize the consistency loss between the input conditional control and extracted condition. A straightforward implementation would be generating images from random noises and then calculating the consistency loss, but such an approach requires storing gradients for multiple sampling timesteps, leading to considerable time and memory costs. To address this, we introduce an efficient reward strategy that deliberately disturbs the input images by adding noise, and then uses the single-step denoised images for reward fine-tuning. This avoids the extensive costs associated with image sampling, allowing for more efficient reward fine-tuning. Extensive experiments show that ControlNet++ significantly improves controllability under various conditional controls. For example, it achieves improvements over ControlNet by 11.1% mIoU, 13.4% SSIM, and 7.6% RMSE, respectively, for segmentation mask, line-art edge, and depth conditions. All the code, models, demo and organized data have been open sourced on our Github Repo.
@@ 1 Introduction
The emergence and improvements of diffusion models [50, 12, 43], along with the introduction of large-scale image-text datasets [49, 48], has catalyzed significant strides in text-to-image generation. Nonetheless, as the proverb “an image is worth a thousand words” conveys, it’s challenging to depict an image accurately and in detail through language alone, and this dilemma also perplexes existing text-to-image diffusion models [43, 46]. To this end, many studies focus on incorporating conditional controls such as segmentation mask into text-to-image diffusion models [63, 30, 62, 37, 22]. Despite the diversity in these methods, the core objective remains to facilitate more accurate and controllable image generation with explicit image-based conditional controls.
Achieving controllable image generation could involve retraining diffusion models from scratch [43, 37], but this comes with high computational demands and a scarcity of large public datasets [63]. In light of this, a feasible strategy is fine-tuning pre-trained text-to-image models [23, 61] or introducing trainable modules [63, 30, 62] like ControlNet [63]. However, despite these studies have explored the feasibility of controllability [63, 30, 62] in text-to-image diffusion models and expanded various applications [23, 37, 22], a significant gap remains in achieving precise and fine-grained control. As shown in Fig. 1, existing methods of controllable generation (e.g., ControlNet [63] and T2I-Adapter [30]) still struggle to accurately generate images that are consistent with the input image condition. For example, T2I-Adapter-SDXL consistently produced incorrect wrinkles on the forehead in all generated images, while ControlNet v1.1 introduced many wrong details. Regrettably, current efforts lack specific methods for improving controllability, which impedes progress in this research area.
@@
  • Consistency Reward Feedback: We show that pre-trained discriminative models can serve as powerful visual reward models to improve the controllability of controllable diffusion models in a cycle-consistency manner.
@@ 2.2 Controllable Text-to-Image Diffusion Models
To achieve conditional control in pre-trained text-to-image diffusion models, ControlNet [63] and T2I-Adapter [30] introduce additional trainable modules for guided image generation. Furthermore, recent research employs various prompt engineering [27, 64, 61] and cross-attention constraints [58, 6, 23] for a more regulated generation. Some methods also explore multi-condition or multi-modal generation within a single diffusion model [37, 21, 65] or focus on the instance-based controllable generation [54, 69]. However, despite these methods exploring feasibility and applications, there still lacks a clear approach to enhance controllability under various controls. Furthermore, existing works implicitly learn controllability by the denoising process of diffusion models, while our ControlNet++ achieves this in an explicit cycle-consistency manner, as shown in Fig. 3.
@@ 3.1 Preliminary
where $\epsilon_{\theta}$ refers to the predicted noise at timestep $t$ by U-Net [44] with parameters $\theta$, and $\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ is the variance of posterior Gaussian distribution $p_{\theta}(x_{0})$.
@@ 3.2 Reward Controllability with Consistency Feedback
where $\lambda$ is a hyper-parameter to adjust the weight of the reward loss. Through this approach, the consistency loss can guide the diffusion model on how to sample at different timesteps to obtain images more consistent with the input controls, thereby enhancing controllability. Nonetheless, directly applying such reward consistency still poses challenges in efficiency in real-world settings.
@@ 3.3 Efficient Reward Fine-tuning
$x_{0}\approx x^{\prime}_{0}=\frac{x^{\prime}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x^{\prime}_{t},c_{v},c_{t},t\right)}{\sqrt{\alpha_{t}}},$ (7)
@@
During the reward fine-tuning phases, we freeze the pre-trained discriminative reward model and text-to-image model, and only update the ControlNet following its original implementation, which ensures the generative capabilities are not compromised. We also observe that using only the reward loss will lead to image distortion, aligning with the conclusions drawn in previous studies [60].
@@ 4.1 Experimental Setup
Given that existing text-image paired datasets for generative models are unable to provide accurate conditional control data pairs [49, 48], such as image-segmentation pairs, we endeavor to select specific datasets for different tasks that can offer more precise image-label data pairs. More specifically, ADE20K [67, 68] and COCOStuff [4] are used for the segmentation mask condition following ControlNet [63]. For the canny edge map, hed edge map, lineart map, and depth map condition, we utilize the MultiGen-20M dataset proposed by UniControl [37], which is a subset of LAION-Aesthetics [48]. For the datasets without text caption such as ADE20K, we utilize MiniGPT-4 [70] to generate the image caption with the instruction “Please briefly describe this image in one sentence”. The training and inference resolution is 512$\times$512 for all datasets and methods. Details are provided in the supplementary material.
@@
We train ControlNet++ on the training set of each corresponding dataset and evaluate all methods on the validation dataset. All the experiments are evaluated under 512$\times$512 resolution for fair comparison. For each condition, we evaluate the controllability by measuring the similarity between the input conditions and the extracted conditions from generated images of diffusion models. For semantic segmentation and depth map controls, we use mIoU and RMSE as evaluation metrics respectively, which is a common practice in related research fields. For the edge task, we use F1-Score for hard edges (canny edge) because it can be regarded as a binary classification problem of 0 (non-edge) and 1 (edge) and has a serious long-tail distribution, following the standard evaluation in edge detection [59]. The threshold used for evaluation is (100, 200) for OpenCV, and (0.1, 0.2) for Kornia implementation. The SSIM metric is used for the soft edges conditional controls (i.e., hed edge & lineart edge) following previous works [65]. For ControlNet++, we use the UniPC [66] sampler with 20 denoising steps to generate images with the original text prompt following ControlNet v1.1 [63], without any negative prompts. For other methods beyond ControlNet and ours, we utilized their open-source code to generate images and conducted fair evaluations under the same data, without changing their inference configures such as the number of inference steps or denoising sampler.
@@
Our evaluation primarily focuses on T2I-Adapter [30], ControlNet v1.1 [63], GLIGEN [27], Uni-ControlNet [65], and UniControl [37], as these methods are pioneering in the realm of controllable text-to-image diffusion models and offer public model weights for various image conditions. To ensure fairness of evaluation, all methods use the same image conditions and text prompts. While most methods employ the user-friendly SD1.5 as their text-to-image model for controllable generation, we have observed that recently there are a few models based on SDXL [35]. Therefore, we also report the controllability results for ControlNet-SDXL and T2I-Adapter-SDXL. Please note that ControlNet-SDXL mentioned here is not an officially released model as in ControlNet [63].
Table 1: Controllability comparison with state-of-the-art methods under different conditional controls and datasets. $\uparrow$ denotes higher result is better, while $\downarrow$ means lower is better. ControlNet++ achieves significant controllability improvements. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images in png format and report the average result to reduce random errors.
@@ 4.2 Experimental Results
The experimental results are shown in Tab. 1, which can be summarized as the following observations: (1) Existing methods still underperform in terms of controllability, struggling to achieve precise controlled generation. For instance, current methods (i.e., ControlNet) achieve only a 32.55 mIoU for images generated under the condition of segmentation masks, which is far from its performance on real datasets with a 50.7 mIoU, under the same evaluation from Mask2Former segmentation model [8]. (2) Our ControlNet++ significantly outperforms existing works in terms of controllability across various conditional controls. For example, it achieves 11.1% RMSE improvements against previous state-of-the-art methods for the depth map condition; (3) For controllable diffusion models, the strength of the text-to-image backbone does not affect its controllability. As shown in the table, although SDXL-based [35] ControlNet and T2I-Adapter have better controllability on some specific tasks, the improvement is not large and is not significantly better than the counterparts with SD 1.5 [43].
Table 2: FID ($\downarrow$) comparison with state-of-the-art methods under different conditional controls and datasets. All the results are conducted on 512$\times$512 image resolution with Clean-FID implementation [33] for fair comparisons. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images in png format and report the average result to reduce random errors.
@@
Table 3: CLIP-score ($\uparrow$) comparison with state-of-the-art methods under different conditional controls and datasets. ‘-’ indicates that the method does not provide a public model for testing. We generate four groups of images in png format and report the average result to reduce random errors.
@@
To verify whether improving controllability leads to a decline in image quality, we reported the FID (Fréchet Inception Distance) metrics of different methods under various conditional generation tasks in Tab. 2. We discovered that, compared to existing methods, ControlNet++ generally exhibits superior FID values in most cases, indicating that our approach, while enhancing the controllability of conditional controls, does not result in a decrease in image quality. This can also be observed in Fig. 6. We provide more visual examples in the supplementary material.
@@
To further validate our improvements in controllability and their impact, we use the generated images along with real human-annotated labels to create a new dataset for training discriminative models from scratch. Please note that the only difference from the original dataset used to train the discriminative model is that we have replaced the images with those generated by the controllable diffusion model while keeping the labels unchanged. If the generative model exhibits good controllability, the quality of the constructed dataset will be higher, thereby enabling to train a stronger model.
@@
Qualitative Comparison.
Figs. 6 and 7 provide a qualitative comparison between our ControlNet++ and previous state-of-the-art methods across different conditional controls. When given the same input text prompts and image-based conditional controls, we observe that existing methods often generate areas
Figure 7: Comparison on Line-Art Edge.
inconsistent with the image conditions. For instance, in the segmentation mask generation task, other methods often produce extraneous picture frames on the walls, resulting in a mismatch between the segmentation masks extracted from the generated image and the inputs. A similar situation occurs under depth conditions, where other methods fail to accurately represent the depth of different fingers. In contrast, images generated by ControlNet++ maintain good consistency with the input depth map.
@@ 4.3 Ablation Study
In Fig. 8, we find that maintaining the original diffusion training process is crucial for preserving the quality and controllability of generated images. Relying solely on pixel-level consistency loss leads to severe image distortion, whereas training the model with both this loss and the diffusion training loss can enhance controllability without affecting image quality.
Generalizability of Efficient Reward Fine-tuning.
Although the reward fine-tuning is used in a small subset of timesteps, it updates all the parameters of the ControlNet and therefore helps more timesteps to improve controllability during sampling. To prove this, we divide the sampling process into two parts: the unoptimized timesteps $[T,t_{thre}]$, and the optimized timesteps $[t_{thre},1]$ and use ControlNet and our model for inference crossly, with 20-step sampling following ControlNet. Table 5 shows that our reward finetuning performed on a small number of timesteps $[t_{thre},1]$ can be generalized to larger timesteps $[T,t_{thre}]$.
Table 4: The impact of efficient reward fine-tuning on different timesteps.
Unoptimized
$[T,t_{thre}]$
Optimized
$[t_{thre},1]$
ADE20K
mIoU ($\uparrow$)
ControlNet ControlNet 32.55
ControlNet Ours 38.03
Ours ControlNet 41.46
Ours Ours 43.64
Table 5: Stronger reward model (UperNet-R50) leads to better controllability than the weaker reward model (DeepLabv3-MBv2).
[…]
- - 32.55
@@
5 Discussion
How to make Hed/LineArt Edge extraction methods differentiable?
The Hed and LineArt Edge extraction models are neural networks without non-differentiable operations. Differentiability can be achieved by simply modifying forward code.
Some conditions (e.g., Box/Sketch/Pose) are not available.
Our reward fine-tuning leverages a pre-trained ControlNet and a differentiable reward model. Currently, pre-trained ControlNet for object bounding boxes and differentiable reward models for sketches are lacking. In existing pose models, there are non-differentiable operations such as the NMS and keypoints grouping. We leave the question of how to extend consistency reward to more conditions to future work.
Influence of Text Prompt.
We discuss how different types of text prompts (No Prompt, Conflicting Prompt, and Perfect Prompt) affect the final results. As shown in Fig. 9, when the text prompt is empty or there is a semantic conflict with the image conditional control, ControlNet often struggles to generate accurate content. In contrast, our ControlNet++ manages to generate images that comply with the input conditional controls under various text prompt scenarios.